# Shortcomings of LLMs for Low-Resource Translation: Retrieval and Understanding are Both the Problem

### v2

@@ 2 LLMs for Machine Translation
Studies to date Elsner and Needle (2023); Reid et al. (2024); Zhang et al. (2024) experiment with four dimensions of variability: source language, LLM, type(s) of information retrieved, and retrieval method. Since the source languages in these studies have relatively little presence in public corpora or on the web, differing results across LLMs can tentatively be attributed to differences in their in-context learning and instruction following abilities.
All studies find that word-level translations are helpful additions to prompts. Zhang et al. (2024) and Tanzer et al. (2024) also add sentence pairs from a parallel corpus, while Elsner and Needle (2023) add usage examples from a dictionary. Each improve results, although to a lesser degree. Elsner and Needle (2023) and Zhang et al. (2024) experiment with small fixed “grammar lesson” passages to provide explicit syntactic instruction, but find these ineffective. Tanzer et al. (2024) uses passages retrieved from a grammar book, also with relatively disappointing results. Reid et al. (2024) use the entire grammar book and a very long-context model to obtain better translations, but without exploring the role explicit grammar instruction actually plays in doing so.
@@ 3.1.1 Morphological Segmentation
S. Quechua is primarily agglutinating, i.e., much of the morphology may be described in terms of isomorphic form-meaning relationships, morphemes generally maintain a consistent form regardless of their phonological environment, and morpheme boundaries tend to be transparent. In contrast, morpheme segmentation in Spanish may be rendered opaque due to its fusional morphology and widespread use of conditional allomorphy.
@@ 4.2.2 Grammar Descriptions (grammar)
We also experiment with the inclusion of grammar lessons found in student-facing pedagogical materials, retrieving grammatical explanations relevant to each source sentence from a PDF document developed for students and teachers of S. Quechua. Pinto Tapia et al. (2005). The document is organized into short sections (1-3 sentences, plus paradigm tables or usage examples) that describe the particular grammatical concept associated with an affix in Quechua. For each source sentence, we retrieve sections associated with any affix listed in the document that is an exact match of a morpheme and include this in prompts using contextual information from the grammar. This improves on the methods described in Tanzer et al. (2024), who use LCS-based retrieval over an entire textbook, and Elsner and Needle (2023); Zhang et al. (2024), whose grammatical description remains consistent across prompts regardless of the source text being translated.
@@ 5.4 Mistranslations
As reported in previous work, adding morpheme- and word-level translations to the prompt greatly reduces the rate of this kind of response. GPT4o also produces drastically fewer mistranslations compared to its predecessor. However, it is notable that both models produce at least one mistranslation for each prompt type. In general, complete mistranslations are in fluent Spanish and contain no overt indications that something has been misrepresented. We return to the ethical implications of these errors in the Discussion.
@@ 7 Conclusion
For the majority of the world’s languages and their speakers, powering and supplying LLMs with enough pretraining data to overcome their limitations is not feasible. We therefore offer the following suggestions to those looking to develop low-resource LLM-MT: (1), improve data structures and methods for interacting with a language-specific database for retrieval-aided generation. (2), continue analysis of the mechanisms driving in-context learning in LLMs, for example by comparing ICL to the effects of finetuning Dai et al. (2023), (3) experiment with prompt structures and techniques, for example by altering the order information Liu et al. (2024) or by iteratively or prompting the model to guide its reasoning towards a suitable translation Wang et al. (2022).
@@ 8 Limitations
We acknowledge, as well, limitations that arise from the size of our dataset and database and the methods used to curate them. The 50 sentence pairs we use were selected to highlight a range of specific grammatical phenomena, not all of which were well represented in our database, and differ in their structural complexity. We are grateful for the guidance provided by the Quechua instructor whose lessons were a source for such examples and proofread the sentences before their inclusion in our experiments, but limited as well by our status as non-native speakers. Human evaluation of model outputs was partially conducted using machine-translated English texts as references, but all annotations were inspected by the Spanish- and Quechua-speaking author who removed a small number of evaluations that reflected linguistic discrepancies between Quechua, Spanish, and English or inaccuracies in the machine translated English.
@@ 9 Ethics Statement
We wish to acknowledge the delicate issue of academic extractiveness and its harmful impact on Indigenous and minority language communities and speakers. We are also aware of some of the controversial ideologies and policies associated with Qheswa Simi Hamut’ana Kuraq Suntur, the government-afilliated institution who published the dictionary we use in this study, and the potentially negative effects of government-sponsored linguistic standardization more broadly (see, e.g., Coronel Molina (2008) for an analysis of the effects of the institution’s ideologies on revitalization efforts in Peru). We do not endorse such policies, and have sought to avoid representing the diversity of Southern Quechua speaking communities as a monolith. Instead, we hope our continued efforts to improve methods for low-resource translation will empower speakers of Southern Quechua and other Indigenous and minority languages to develop language technologies capable of representing their own community’s unique language variety to serve the unique needs of its speakers.

### v3

@@ 2 LLMs for Machine Translation
Studies to date Elsner and Needle (2023); Reid et al. (2024); Zhang et al. (2024) experiment with four dimensions of variability: source language, LLM, type(s) of information retrieved, and retrieval method. Since the source languages in these studies have relatively little presence in public corpora or on the web, differing results across LLMs can tentatively be attributed to differences in their in-context learning and instruction-following abilities.
All studies find that word-level translations are helpful additions to prompts. Zhang et al. (2024) and Tanzer et al. (2024) also add sentence pairs from a parallel corpus, while Elsner and Needle (2023) add usage examples from a dictionary. Each improves results, although to a lesser degree. Elsner and Needle (2023) and Zhang et al. (2024) experiment with small fixed “grammar lesson” passages to provide explicit syntactic instruction, but find these ineffective. Tanzer et al. (2024) uses passages retrieved from a grammar book, also with relatively disappointing results. Reid et al. (2024) use the entire grammar book and a very long-context model to obtain better translations, but without exploring the role explicit grammar instruction actually plays in doing so.
@@ 3.1.1 Morphological Segmentation
S. Quechua is primarily agglutinating, i.e., much of the morphology may be described in terms of isomorphic form-meaning relationships, morphemes generally maintain a consistent form regardless of their phonological environment, and morpheme boundaries tend to be transparent. In contrast, morpheme segmentation in Spanish may be rendered opaque due to its fusional morphology and widespread use of conditioned allomorphy.
@@ 4.2.2 Grammar Descriptions (grammar)
We also experiment with the inclusion of grammar lessons found in student-facing pedagogical materials, retrieving grammatical explanations relevant to each source sentence from a PDF document developed for students and teachers of S. Quechua Pinto Tapia et al. (2005). The document is organized into short sections (1-3 sentences, plus paradigm tables or usage examples) that describe the particular grammatical concept associated with an affix in Quechua. For each source sentence, we retrieve sections associated with any affix listed in the document that is an exact match of a morpheme and include this in prompts using contextual information from the grammar. This improves on the methods described in Tanzer et al. (2024), who use LCS-based retrieval over an entire textbook, and Elsner and Needle 2023 and Zhang et al. 2024, whose grammatical description remains consistent across prompts regardless of the source text being translated.
@@ 5.4 Mistranslations
As reported in previous work, adding morpheme- and word-level translations to the prompt greatly reduces the rate of this kind of response. GPT-4o also produces drastically fewer mistranslations compared to its predecessor. However, it is notable that both models produce at least one mistranslation for each prompt type. In general, complete mistranslations are in fluent Spanish and contain no overt indications that something has been misrepresented. We return to the ethical implications of these errors in the Discussion.
@@ 7 Conclusion
For the majority of the world’s languages and their speakers, powering and supplying LLMs with enough pretraining data to overcome their limitations is not feasible. We therefore offer the following suggestions to those looking to develop low-resource LLM-MT: (1) improve data structures and methods for interacting with a language-specific database for retrieval-aided generation, (2) continue analysis of the mechanisms driving in-context learning in LLMs, for example by comparing ICL to the effects of finetuning Dai et al. (2023), and (3) experiment with prompt structures and techniques, for example by altering the order of information Liu et al. (2024) or by iteratively prompting the model to guide its reasoning towards a suitable translation Wang et al. (2022).
@@ 8 Limitations
We acknowledge, as well, limitations that arise from the size of our dataset and database and the methods used to curate them. The 50 sentence pairs we use were selected to highlight a range of specific grammatical phenomena, not all of which were well represented in our database, and differ in their structural complexity. We are grateful for the guidance provided by the Quechua instructor whose lessons were a source for such examples and proofread the sentences before their inclusion in our experiments, but are limited by our status as non-native speakers. Human evaluation of model outputs was partially conducted using machine-translated English texts as references, but all annotations were inspected by the Spanish- and Quechua-speaking author who removed a small number of evaluations that reflected linguistic discrepancies between Quechua, Spanish, and English or inaccuracies in the machine-translated English.
@@ 9 Ethics Statement
We wish to acknowledge the delicate issue of academic extractiveness and its harmful impact on Indigenous and minority language communities and speakers. We are also aware of some of the controversial ideologies and policies associated with Qheswa Simi Hamut’ana Kuraq Suntur, the government-afilliated institution who published the dictionary we use in this study, and the potentially negative effects of government-sponsored linguistic standardization more broadly (see, e.g., Coronel Molina (2008) for an analysis of the effects of the institution’s ideologies on revitalization efforts in Peru). We do not endorse such policies, and have sought to avoid representing the diversity of Southern Quechua-speaking communities as a monolith. Instead, we hope our continued efforts to improve methods for low-resource translation will empower speakers of Southern Quechua and other Indigenous and minority languages to develop language technologies capable of representing their own community’s unique language variety to serve the unique needs of its speakers.